OpenStudy (anonymous):

Sketch the right triangle and find the length of the side not given. If necessary, approximate the length to the nearest thousandth. leg = 12, hypotenuse = 20 A. 19 B. 14 C. 20 D. 16

OpenStudy (anonymous):

answer is D..?

OpenStudy (anonymous):

Use the pythagorean theorem to find legs of a right triangle. (a^2+b^2 = c^2) where c is your hypotenuse. Therefore: 12^2 + b^2 = 20^2. Solve for B. = 16. D